OpenStudy (anonymous):

A boat traveled downstream a distance of 56 mi abd tgeb cane right back. If the speed of the current was 5 mph and then total trip took 4 hours and 40 minutes, find the average speed of the boat relative to the water

OpenStudy (p0sitr0n):

time: 14/3 hours 56/(5+x) + 56/(5-x) = 14/3 56 ( 1/5+x + 1/5-x)=14/3 56/(10/25-x^2)=14/3 12 = 10/25-x^2 25-x^2 = 10/12 x=4.9 mph